# Auslander-Reiten translations in the monomorphism categories of exact categories

### v1

@@ 3.1. Exact structures on morphism categories
[HE, Proposition 3.2] Assume $\mathcal{C}$ is idempotent-complete and admits pseudo-kernels. Then ${\mathcal{E}}^{\rm cw}$ defines an exact structure on the additive category ${{\mathcal{H}}}({\mathcal{C}})$.
When ${\mathcal{C}}$ satisfies the above proposition, we denote the exact category $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm cw})$ by ${{\mathcal{H}}}^{\rm cw}({\mathcal{C}})$.
If ${\mathcal{C}}$ is moreover an exact category, then ${\mathcal{H}}({\mathcal{C}})$ enjoys a richer exact structure, which is induced naturally from the exact structure of ${\mathcal{C}}$.
Definition 3.4 (The deg-exact structure).
Let ${\mathcal{E}}^{\rm deg}$ be the class of all pairs of composable morphisms
such that the induced composable morphisms $X_{i}\stackrel{{\scriptstyle\phi_{i}}}{{\rightarrow}}Y_{i}\stackrel{{\scriptstyle\psi_{i}}}{{\rightarrow}}Z_{i}$ exact in ${\mathcal{C}}$ for $i=1,2$.
When ${\mathcal{C}}$ is exact, there are two exact structures on ${\mathcal{H}}({\mathcal{C}})$, namely ${\mathcal{E}}^{\rm cw}$ and ${\mathcal{E}}^{\rm deg}$. It is clear that ${\mathcal{E}}^{\rm cw}\subseteq{\mathcal{E}}^{\rm deg}$. So any cw-exact sequence is still exact in $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$. Intuitively, ${\mathcal{E}}^{\rm deg}$ is the more natural exact structure because when ${\mathcal{C}}\subseteq\Lambda\textendash\operatorname{mod}$, $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$ is an exact subcategory of ${\mathcal{H}}(\Lambda)\cong T_{2}(\Lambda)\textendash\operatorname{mod}$. Hence, in the following, when we refer to the exact category ${{\mathcal{H}}}({\mathcal{C}})$ without emphasizing the exact structure, we mean $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$. Although the exact structure ${\mathcal{E}}^{\rm cw}$ is much smaller than ${\mathcal{E}}^{\rm deg}$ in general, it is surprising that when $({\mathcal{H}}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$ has almost split sequences, most of the almost split sequences are cw-exact. Hence, when it comes to characterizing almost split sequences, ${\mathcal{H}}^{\rm cw}({\mathcal{C}})$ becomes a convenient tool.
3.2. Functors on morphisms categories
@@
Let $F:{\mathcal{C}}\to{\mathcal{D}}$ be an additive functor. If $F$ is an equivalence, then so is ${\mathcal{H}}(F)$.
@@
It suffices to prove ${\mathcal{H}}(F)$ is dense and fully-faithful. For the density of ${\mathcal{H}}(F)$, let $\left(\begin{smallmatrix}A^{\prime}_{1}\\ A^{\prime}_{2}\end{smallmatrix}\right)_{g}\in{\mathcal{H}}({\mathcal{D}})$. By the density of $F$, there are $A_{1},A_{2}\in{\mathcal{C}}$, such that $F(A_{1})\cong A^{\prime}_{1},F(A_{2})\cong A^{\prime}_{2}$. Since $F$ is full, there is a homomorphism $f:A_{1}\to A_{2}$ such that $F(f)\cong g.$ So ${\mathcal{H}}(F)$ is dense. For the fullness of ${\mathcal{H}}(F)$, let $\left(\begin{smallmatrix}A_{1}\\ A_{2}\end{smallmatrix}\right)_{f}$ and $\left(\begin{smallmatrix}B_{1}\\ B_{2}\end{smallmatrix}\right)_{g}$ be objects in ${\mathcal{C}}$ and $(\alpha^{\prime},\beta^{\prime}):Ff\to Fg$ be a homomorphism in ${\mathcal{H}}({\mathcal{D}})$, i.e. $Fg\alpha^{\prime}=\beta^{\prime}Ff$. By the fullness of $F$, there are homomorphisms $\alpha:A_{1}\to A_{2}$ and $\beta:B_{1}\to B_{2}$ such that $F\alpha=\alpha^{\prime}$ and $F\beta=\beta^{\prime}$. Hence $FgF\alpha=F\beta Ff$. Since $F$ is a faithful functor, we have $g\alpha=\beta f$. That is to say, $(\alpha,\beta):f\to g$ is a homomophism in ${\mathcal{H}}({\mathcal{C}})$ such that $F(\alpha,\beta)=(\alpha^{\prime},\beta^{\prime})$. Hence ${\mathcal{H}}(F)$ is full. It is easy to see that the faithfulness of ${\mathcal{H}}(F)$ is guaranteed by the faithfulness of $F$. This completes the proof. ∎
Let ${\mathcal{C}}$ be an exact category. Define the monomorphism category $\mathcal{S}({\mathcal{C}})$ (resp. epimorphism category $\mathcal{F}({\mathcal{C}})$) as the full subcategory of $(\mathcal{H}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$, consisting of objects which are monomorphisms (resp. epimorphisms).
There are a couple of functors:
@@
Additive functors $\operatorname{\bf Ker}$ and $\operatorname{\bf Cok}$ form a pair of inverse equivalences.
As immediate consequences of the RSS equivalence. We will see that $\operatorname{\bf Ker}$ and $\operatorname{\bf Cok}$ preserve indecomposable objects, projectives, injectives and almost split sequences.
3.3. Almost split sequences in morphism categories
The following result is due to [HE, Proposition 4.6], which asserts the existence of almost split sequences in the morphism categories.
Proposition 3.7.
Let ${\mathcal{C}}$ be an extension-closed $k$-dualizing subvariety of $\Lambda\textendash\operatorname{mod}$. Then the canonical exact category $(\mathcal{H}({\mathcal{C}}),{\mathcal{E}}^{\rm deg})$ admits almost split sequences.
Definition 3.8.
An indecomposable object in ${\mathcal{H}}({\mathcal{C}})$ is said to be of type $(a)$ (resp. $(b)$, or $(c)$) provided it is isomorphic to $(0\rightarrow M)$ (resp. $(M\stackrel{{\scriptstyle 1}}{{\rightarrow}}M)$, or $(M\rightarrow 0)$) for some $\Lambda$-module $M$. Furthermore, an indecomposable object is said to be of type $(d)$ if it is isomorphic to $(P\stackrel{{\scriptstyle f}}{{\rightarrow}}Q)$ where $P,Q$ are projective $\Lambda$-modules.
The following characterization of almost split sequences can be obtained from the definitions.
Lemma 3.9.
[HE, Lemma 4.5] Assume $\delta:0\rightarrow A\stackrel{{\scriptstyle f}}{{\rightarrow}}B\stackrel{{\scriptstyle g}}{{\rightarrow}}C\rightarrow 0$ is an almost split sequence in ${\mathcal{C}}$. Then
  • $(1)$
    The almost split sequence in ${{\mathcal{H}}}({\mathcal{C}})$ ending at $(0\rightarrow C)$ is of the form
@@ 3.3. Almost split sequences in morphism categories
  • $(2)$
    The almost split sequence in ${{\mathcal{H}}}({\mathcal{C}})$ ending at $(C\stackrel{{\scriptstyle 1}}{{\rightarrow}}C)$ is of the form
@@
The following corollary is an immediate consequence of the arguments above.
Corollary 3.10.
Let ${\mathcal{C}}$ be an extension-closed $k$-dualizing subvariety of $\Lambda\textendash\operatorname{mod}$ and let
[…]

### v2

@@ 7. Monomorphism categories of Frobenius categories
For any object $f\in\overline{{\mathcal{S}}({\mathcal{C}})}$, $f\langle 1\rangle\cong\operatorname{\bf Mimo}(\overline{f}[1])$.
Proof.
It suffices to prove for indecomposable objects $\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)_{f}\in{\mathcal{S}}({\mathcal{C}})_{\mathcal{I}}$. Suppose $i_{B}:B\to I(B)$ is an $\operatorname{Ext}$-injective envelope and $i_{A}=i_{B}\circ f$. Then there is a commutative diagram with exact rows:
[…]
where $f^{\prime}$ is an epimorphism satisfying $\overline{f^{\prime}}\cong\overline{f}[1]$ in $\overline{{\mathcal{C}}}$. Let $\operatorname{\bf Mimo}f^{\prime}=[f^{\prime},e]:A^{\prime}\to B^{\prime}\oplus I^{\prime}$. Then it is easy to check that there is an exact sequence in ${\mathcal{S}}({\mathcal{C}})$ below:
[…]
Hence, in the stable category $\overline{{\mathcal{S}}({\mathcal{C}})}$, we have $f\langle 1\rangle\cong\operatorname{\bf Mimo}f^{\prime}\cong\operatorname{\bf Mimo}(\overline{f}[1])$ by the Remark 5.9. ∎
Throughout the rest of this section, let $\mathcal{V}$} be a dualizing $k$-variety and $\mathcal{C}$} be a functorially finite Frobenius subcategory of ${\mathcal{V}}\textendash\operatorname{mod}$ having enough $\operatorname{Ext}$-projective and $\operatorname{Ext}$-injective objects. Denote by $\tau$ the Auslander-Reiten translation on $\mathcal{C}$}. According to Theorem A and Lemma 7.1, ${\mathcal{S}}({\mathcal{C}})$ is again a Frobenius category having Auslander-Reiten-Serre duality. Denote by $\tau_{\mathcal{S}}$ the Auslander-Reiten translation on ${\mathcal{S}}({\mathcal{C}})$. Notice that triangle functors $\tau$ and $\tau_{\mathcal{S}}$ yield auto-equivalences of the stable categories $\overline{{\mathcal{C}}}$ and $\overline{{\mathcal{S}}({\mathcal{C}})}$ respectively.
To deduce a formula for the triangle functor $\tau_{\mathcal{S}}$, we need the following notation.
Definition 7.3.
Let $\left(\begin{smallmatrix}X\\ Y\end{smallmatrix}\right)_{\overline{f}}\in{\mathcal{H}}(\overline{{\mathcal{C}}})$, which fits into a distinguished triangle $X\stackrel{{\scriptstyle\overline{f}}}{{\to}}Y\stackrel{{\scriptstyle\overline{g}}}{{\to}}Z\stackrel{{\scriptstyle\overline{h}}}{{\to}}X[1]$ in $\overline{{\mathcal{C}}}$, the rotation operation $\operatorname{\bf Rot}\left(\begin{smallmatrix}X\\ Y\end{smallmatrix}\right)_{\overline{f}}$ is defined to be $\left(\begin{smallmatrix}Y\\ Z\end{smallmatrix}\right)_{\overline{g}}$.
We will write $\operatorname{\bf Rot}(\overline{f})$ for convenience. From the axioms of triangulated category, it is clear that $\operatorname{\bf Rot}$ is well-defined on objects of ${\mathcal{H}}(\overline{{\mathcal{C}}})$. However, we warn the reader that in general it is not a functor.
@@
  1. (1)
    For any $\overline{f}\in{{\mathcal{H}}(\overline{{\mathcal{C}}})}$, $\operatorname{\bf Rot}^{3}\overline{f}\cong-\overline{f}[1]$ and $(\operatorname{\bf Rot}\overline{f})[1]=\operatorname{\bf Rot}(\overline{f}[1])$.
  2. (2)
    For any $f\in\overline{{\mathcal{S}}({\mathcal{C}})}$, $\overline{\operatorname{\bf Cok}f}=\operatorname{\bf Rot}\overline{f}$.
@@
The formula $\operatorname{\bf Rot}^{3}\overline{f}\cong-\overline{f}[1]$ follows from the axioms of triangulated category directly. For the formula $\overline{\operatorname{\bf Cok}f}=\operatorname{\bf Rot}\overline{f}$, its proof is essentially the same as Lemma 6.1 in [RS2]. Here we only need to prove $(\operatorname{\bf Rot}\overline{f})[1]=\operatorname{\bf Rot}(\overline{f}[1])$. Let $X\stackrel{{\scriptstyle\overline{f}}}{{\to}}Y\stackrel{{\scriptstyle\overline{g}}}{{\to}}Z\stackrel{{\scriptstyle\overline{h}}}{{\to}}X[1]$ be a distinguished triangle in $\overline{{\mathcal{C}}}$. Then by the axioms of triangulated category, $X[1]\stackrel{{\scriptstyle-\overline{f}[1]}}{{\longrightarrow}}Y[1]\stackrel{{\scriptstyle-\overline{g}[1]}}{{\longrightarrow}}Z[1]\stackrel{{\scriptstyle-\overline{h}[1]}}{{\longrightarrow}}X[2]$ is a distinguished triangle, which is isomorphic to the triangle $X[1]\stackrel{{\scriptstyle\overline{f}[1]}}{{\longrightarrow}}Y[1]\stackrel{{\scriptstyle\overline{g}[1]}}{{\longrightarrow}}Z[1]\stackrel{{\scriptstyle-\overline{h}[1]}}{{\longrightarrow}}X[2]$. Hence by the definition of $\operatorname{\bf Rot}$, we have $(\operatorname{\bf Rot}\overline{f})[1]=\operatorname{\bf Rot}(\overline{f}[1])$. ∎
Recall that the stable quotient $\pi:\overline{{\mathcal{S}}({\mathcal{C}})}\to{\mathcal{H}}(\overline{{\mathcal{C}}})$ is a full and dense functor.
Lemma 7.5.
Let $f\in\overline{{\mathcal{S}}({\mathcal{C}})}$. Then $\overline{\tau_{\mathcal{S}}f}\cong\tau\operatorname{\bf Rot}\overline{f}$ in ${\mathcal{H}}(\overline{{\mathcal{C}}})$.
Proof.
By Theorem A and Proposition 7.4 (2), $\overline{\tau_{\mathcal{S}}f}\cong\overline{\operatorname{\bf Mimo}\tau\underline{\operatorname{\bf Cok}f}}\cong\tau\overline{\operatorname{\bf Cok}f}\cong\tau\operatorname{\bf Rot}\overline{f}$. ∎
7.2. Stably Calabi-Yau Frobenius categories
Recall that on a $k$-linear $\operatorname{Hom}$-finite additive category $\mathcal{A}$}, a right Serre functor is an additive functor $F:{\mathcal{A}}\to{\mathcal{A}}$ together with isomorphisms:
$\eta_{A,B}:\operatorname{Hom}(A,B)\to D\operatorname{Hom}(B,FA)$
for any $A,B\in{\mathcal{A}}$ and natural in $A$ and $B$. A left Serre functor $G$ is an additive functor $G:{\mathcal{A}}\to{\mathcal{A}}$ together with natural isomorphisms
$\xi_{A,B}:\operatorname{Hom}(A,B)\to D\operatorname{Hom}(GB,A).$
Any left or right Serre functor is full-faithful [RV]. A Serre functor is a right Serre functor $\mathbb{S}$} which is dense. It follows that $\mathcal{A}$} admits a Serre functor if and only if it has both a left and a right Serre functor [RV].
A $k$-linear $\operatorname{Hom}$-finite triangulated category $\mathcal{T}$} is called $d$-Calabi-Yau, if $\mathcal{T}$} has a Serre functor ${\mathbb{S}}=[d]$, i.e. there is a bifunctorial isomorphisms $\operatorname{Hom}(A,B)\cong D\operatorname{Hom}(B,A[d])$ for all $A,B\in{\mathcal{T}}$. Due to a well-known result from [RV], for a $k$-linear $\operatorname{Hom}$-finite triangulated category $\mathcal{T}$}, $\mathcal{T}$} has Auslander-Reiten triangles if and only if it admits a Serre functor.
A $k$-linear $\operatorname{Hom}$-finite Frobenius category $\mathcal{C}$} is called stably $d$-Calabi-Yau, if the stable category $\underline{{\mathcal{C}}}$ is a $d$-Calabi-Yau triangulated category. Furthermore for a stably $d$-Calabi-Yau Frobenius category, there is a functorial isomorphism $\tau\cong[d-1]$, where $\tau:\underline{{\mathcal{C}}}\to\underline{{\mathcal{C}}}$ is the Auslander-Reiten translation.
Applying Theorem A for the monomorphism category ${\mathcal{S}}({\mathcal{C}})$, we have the following computation:
$\tau_{\mathcal{S}}f$ $\cong\operatorname{\bf Mimo}\overline{\tau_{\mathcal{S}}f}$ (Lemma 5.10)
$\cong\operatorname{\bf Mimo}\tau\operatorname{\bf Rot}\overline{f}.$ (Lemma 7.5)
Therefore $\tau_{\mathcal{S}}^{2}f\cong\operatorname{\bf Mimo}\tau\operatorname{\bf Rot}\overline{\operatorname{\bf Mimo}\tau\operatorname{\bf Rot}\overline{f}}\cong\operatorname{\bf Mimo}\tau\operatorname{\bf Rot}\tau\operatorname{\bf Rot}\overline{f}$.
So by induction,
$\tau^{6}_{\mathcal{S}}f$ $\cong\operatorname{\bf Mimo}(\tau\operatorname{\bf Rot})^{6}\overline{f}$
$\cong\operatorname{\bf Mimo}([d-1]\operatorname{\bf Rot})^{6}\overline{f}$
$\cong\operatorname{\bf Mimo}\operatorname{\bf Rot}^{6}(\overline{f}[6d-6])$
$\cong\operatorname{\bf Mimo}\overline{f}[6d-4]$
$\cong f\langle 6d-4\rangle.$ (Prop. 7.2)
This proves Theorem B.
Remark 7.6.
Theorem B suggests that the monomorphism category ${\mathcal{S}}({\mathcal{C}})$ of a stably $d$-Calabi-Yau Frobenius category $\mathcal{C}$} has likely a fractional Calabi-Yau-dimension. Indeed, we can also compute $\tau_{\mathcal{S}}f$ as:
$\tau_{\mathcal{S}}f$ $\cong\operatorname{\bf Mimo}\overline{\tau_{\mathcal{S}}f}$ (Lemma 5.10)
$\cong\operatorname{\bf Mimo}\tau\overline{\operatorname{\bf Cok}f}$
$\cong\operatorname{\bf Mimo}(\overline{\operatorname{\bf Cok}f}[d-1])$
$\cong(\operatorname{\bf Mimo}\overline{\operatorname{\bf Cok}f})\langle d-1\rangle$ (Prop.7.2)
$\cong(\operatorname{\bf Mimo}\operatorname{\bf Cok}f)\langle d-1\rangle$
Notice that $\operatorname{\bf Mimo}\operatorname{\bf Cok}\langle d-1\rangle$ is a functor $\overline{{\mathcal{S}}({\mathcal{C}})}\to\overline{{\mathcal{S}}({\mathcal{C}})}$. Although on many examples we can verify that $\operatorname{\bf Mimo}\operatorname{\bf Cok}\langle d\rangle$ is the Serre functor, it is still an open question for us to prove the functorial isomorphism $\tau_{\mathcal{S}}\cong\operatorname{\bf Mimo}\operatorname{\bf Cok}\langle d-1\rangle$.
8. Examples
8.1. The monomorphism power categories
Let $\mathcal{C}$} be an exact category. Inductively, set ${\mathcal{S}}^{0}({\mathcal{C}})={\mathcal{C}}$ and for a natural number $n\geq 1$ define the monomorphism power category ${\mathcal{S}}^{n}({\mathcal{C}}):={\mathcal{S}}({\mathcal{S}}^{n-1}({\mathcal{C}}))$. As an immediate consequence of Lemma 7.1, we have the following:
@@ 8.1. The monomorphism power categories
Let $\mathcal{C}$} be a Frobenius subcategory of $\Lambda\textendash\operatorname{mod}$. Then the monomorphism category ${\mathcal{S}}^{n}({\mathcal{C}})$ is a Frobenius category.
Example 8.2.
Let $A=k[x]/\langle x^{2}\rangle$. We compute the Auslander-Reiten quiver for ${\mathcal{S}}^{2}(A\textendash\operatorname{mod})$.
Let $S$ be the unique simple module of $A$. By the construction of ${\mathcal{S}}^{2}(A\textendash\operatorname{mod})$, we know that ${\mathcal{S}}^{2}(A\textendash\operatorname{mod})=\{\rm{inflations\ in\ }{\mathcal{S}}(A\textendash\operatorname{mod})\}$. For the inflation $\phi:\left(\begin{smallmatrix}X\\ Y\end{smallmatrix}\right)_{f}\to\left(\begin{smallmatrix}X^{\prime}\\ Y^{\prime}\end{smallmatrix}\right)_{f^{\prime}}$ which is indecomposable in ${\mathcal{S}}^{2}(A\textendash\operatorname{mod})$, we simply write it as $\begin{matrix}{\begin{smallmatrix}X\\ Y\end{smallmatrix}}\\ {\begin{smallmatrix}X^{\prime}\\ Y^{\prime}\end{smallmatrix}}\end{matrix}$, if $\phi$ is uniquely determined by $X,Y,X^{\prime},Y^{\prime}$. By theorem $A$, We can compute the Auslander-Reiten quiver for ${\mathcal{S}}^{2}(A\textendash\operatorname{mod})$ as follows:
$\begin{matrix}{\begin{smallmatrix}0\\ S\end{smallmatrix}}\\ {\begin{smallmatrix}0\\ A\end{smallmatrix}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}S\\ S\end{smallmatrix}}\\ {\begin{smallmatrix}S\\ A\end{smallmatrix}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}S\\ A\end{smallmatrix}}\\ {\begin{smallmatrix}S\\ A\end{smallmatrix}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}S\\ A\end{smallmatrix}}\\ {{\begin{matrix}{\begin{smallmatrix}0\\ S\end{smallmatrix}}\oplus{\begin{smallmatrix}A\\ A\end{smallmatrix}}\end{matrix}}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}0\\ 0\end{smallmatrix}}\\ {\begin{smallmatrix}S\\ S\end{smallmatrix}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}0\\ S\end{smallmatrix}}\\ {{\begin{matrix}{\begin{smallmatrix}S\\ S\end{smallmatrix}}\oplus{\begin{smallmatrix}0\\ A\end{smallmatrix}}\end{matrix}}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}0\\ S\end{smallmatrix}}\\ {\begin{smallmatrix}0\\ A\end{smallmatrix}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}S\\ S\end{smallmatrix}}\\ {\begin{smallmatrix}S\\ S\end{smallmatrix}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}0\\ A\end{smallmatrix}}\\ {\begin{smallmatrix}0\\ A\end{smallmatrix}}\end{matrix}$$\begin{smallmatrix}{\begin{smallmatrix}0\\ 0\end{smallmatrix}}\\ {\begin{smallmatrix}0\\ S\end{smallmatrix}}\end{smallmatrix}$$\begin{matrix}{\begin{smallmatrix}S\\ A\end{smallmatrix}}\\ {{\begin{matrix}{\begin{smallmatrix}A\\ A\end{smallmatrix}}\oplus{\begin{smallmatrix}0\\ A\end{smallmatrix}}\end{matrix}}}\end{matrix}$$\begin{smallmatrix}\begin{smallmatrix}S\\ S\end{smallmatrix}\\ \begin{smallmatrix}S\\ S\end{smallmatrix}\end{smallmatrix}$$\begin{matrix}{\begin{smallmatrix}0\\ 0\end{smallmatrix}}\\ {\begin{smallmatrix}S\\ A\end{smallmatrix}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}S\\ S\end{smallmatrix}}\\ {\begin{smallmatrix}A\\ A\end{smallmatrix}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}0\\ S\end{smallmatrix}}\\ {\begin{smallmatrix}0\\ S\end{smallmatrix}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}0\\ 0\end{smallmatrix}}\\ {\begin{smallmatrix}S\\ A\end{smallmatrix}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}0\\ 0\end{smallmatrix}}\\ {\begin{smallmatrix}0\\ A\end{smallmatrix}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}A\\ A\end{smallmatrix}}\\ {\begin{smallmatrix}A\\ A\end{smallmatrix}}\end{matrix}$$\begin{matrix}{\begin{smallmatrix}0\\ 0\end{smallmatrix}}\\ {\begin{smallmatrix}A\\ A\end{smallmatrix}}\end{matrix}$}
A finite dimensional Iwanaga-Gorenstein algebra $A$ is called stably $d$-Calabi-Yau, if the category of Gorenstein projective $A$-modules is a stably $d$-Calabi-Yau Frobenius category. A finite dimensional algebra $A$ is called $d$-Calabi-Yau tilted if it is isomorphic to the endomorphism algebra of some $d$-cluster tilting object in a $d$-Calabi-Yau triangulated category. Keller and Reiten showed that if $A$ is a 2-Calabi-Yau tilted algebras, then $A$ is 1-Iwanaga-Gorenstein and stably 3-Calabi-Yau [KR]. Hence the subcategory of Gorenstein projective $A$-modules ${\rm Gproj}(A)$ is a functorially finite and stably 3-Calabi-Yau Frobenius subcategory. Let $\Lambda=\begin{bmatrix}A&A\\ 0&A\end{bmatrix}$ be the upper triangular matrix algebra. Since $\Lambda\textendash\operatorname{mod}\cong{\mathcal{H}}(A)$, every $\Lambda$-module is an $A$-homomorphism $\left(\begin{smallmatrix}X\\ Y\end{smallmatrix}\right)_{f}$ and ${\rm Gproj}(\Lambda)\cong{\mathcal{S}}({\rm Gproj}(A))$ (Theorem 1.2). By Remark 7.6, $\tau_{Gp(\Lambda)}\left(\begin{smallmatrix}X\\ Y\end{smallmatrix}\right)_{f}\cong\operatorname{\bf Mimo}\operatorname{\bf Cok}\left(\begin{smallmatrix}X\\ Y\end{smallmatrix}\right)_{f}\langle 2\rangle.$
Example 8.3.
Let $A$ be the following bounded quiver algebra with relation $\beta\delta,\delta\gamma,\gamma\beta$.
@@
The algebra $A$ is cluster-tilted, and hence $2$-Calabi–Yau tilted. Therefore ${\rm Gproj}(A)$ is a $3$-Calabi-Yau Frobenius category.
@@
Figure 3. Auslander-Reiten quiver of $A\textendash\operatorname{mod}$
In Figure 3, $P_{2}=\begin{smallmatrix}&2\\ 1&&3\end{smallmatrix}$, $P_{3}=\begin{smallmatrix}3\\ 4\end{smallmatrix}$, $P_{4}=\begin{smallmatrix}4\\ 2\\ 1\end{smallmatrix}$, $rP_{4}=\begin{smallmatrix}2\\ 1\end{smallmatrix}$, $I_{2}=\begin{smallmatrix}4\\ 2\end{smallmatrix}$, $I_{3}=\begin{smallmatrix}2\\ 3\end{smallmatrix}$ and ${\rm Gproj}(A)$ are marked red. The monomorphism category ${\mathcal{S}}({\rm Gproj}(A))$ contains $17$ indecomposable objects.
$\begin{smallmatrix}3\\ P_{2}\end{smallmatrix}$$\begin{smallmatrix}rP_{4}\\ rP_{4}\end{smallmatrix}$$\begin{smallmatrix}0\\ 4\end{smallmatrix}$$\begin{smallmatrix}4\\ P_{3}\end{smallmatrix}$$\begin{smallmatrix}3\\ 3\end{smallmatrix}$$\begin{smallmatrix}0\\ rP_{4}\end{smallmatrix}$$\begin{smallmatrix}rP_{4}\\ P_{4}\end{smallmatrix}$$\begin{smallmatrix}4\\ 4\end{smallmatrix}$$\begin{smallmatrix}0\\ 3\end{smallmatrix}$$\begin{smallmatrix}3\\ P_{2}\end{smallmatrix}$$\begin{smallmatrix}P_{2}\\ P_{2}\end{smallmatrix}$$\begin{smallmatrix}0\\ P_{3}\end{smallmatrix}$$\begin{smallmatrix}P_{3}\\ P_{3}\end{smallmatrix}$$\begin{smallmatrix}0\\ P_{4}\end{smallmatrix}$$\begin{smallmatrix}P_{4}\\ P_{4}\end{smallmatrix}$$\begin{smallmatrix}0\\ P_{2}\end{smallmatrix}$$\begin{smallmatrix}1\\ 1\end{smallmatrix}$$\begin{smallmatrix}1\\ 1\end{smallmatrix}$$\begin{smallmatrix}0\\ 1\end{smallmatrix}$
Figure 4. Auslander-Reiten quiver of ${\mathcal{S}}({\rm Gproj}(A))$
Acknowledgments: We would like to thank Rasool Hafezi, Sondre Kvamme, Claus Michael Ringel and Markus Schmidmeier for helpful discussions and comments.